1)simplify 4squareroot16x4y10
2)simplify squareroot60xy7z4

Answer by checkley75(3666)

1) 4SQRT16X^4Y^10=4*4X^2Y^5=16X^2Y^5 ANSWER.
2) SQRT60XY^7Z^4=2Y^3Z^2SQRT15XY ANSWER.